# Velocity-resolved Reverberation Mapping of Changing-look Active Galactic Nucleus NGC 4151 during Outburst Stage. II. Four Season Observation Results

### v1

@@ 1 Introduction
Over the past few decades, extensive observational and theoretical studies have revealed the basic picture of AGN structure: the central SMBH produces most of the observed continuum radiation through an accretion disk, which is surrounded by the broad-line region (BLR) composed of rapidly orbiting ionized gas emitting broad emission lines, a dust torus that can obscure the inner regions depending on the viewing angle, and the narrow-line region extending to larger scales and emitting narrow emission lines (Antonucci, 1993; Urry & Padovani, 1995). Despite the huge success of the unified model in explaining the diverse observational features of AGNs, there remain many mysteries regarding the detailed physical processes and evolution of these structures, such as the observed accretion disk sizes often being larger than theoretically expected, the complex geometry and kinematics of the BLR, and the influence of radiation pressure on the internal physical properties of AGNs (e.g., Gaskell & Harrington, 2018; Pancoast et al., 2018; Villafaña et al., 2022; Chen et al., 2023; Liu et al., 2024; Li et al., 2024; Zhou et al., 2024).
@@
Additionally, some high-quality RM data allow for a detailed analysis of the emission line response at different line-of-sight velocities, enabling the inference of the geometry and kinematic properties of the BLR. This approach is known as the “velocity-resolved” RM method. These studies have revealed that, in addition to Keplerian motion, the BLR often exhibits complex kinematic features such as inflows or outflows (Denney et al., 2009; Bentz et al., 2010; Grier et al., 2013; Pancoast et al., 2014; Du et al., 2016; Lu et al., 2019; Li et al., 2021). Intriguingly, RM observations of the same AGN at different durations sometimes show significant changes in the kinematic features of the BLR (Grier et al., 2017; Lu et al., 2022; Chen et al., 2023). It is generally believed that changes in BLR kinematics may be related to variations in the accretion rate (Proga et al., 2000; Kollatschny & Zetzl, 2013), where higher accretion rates could lead to stronger radiation pressure driving more outflowing gas. However, Du et al. (2016) found no significant correlation between BLR kinematics and accretion rates through a sample of high accretion rate AGNs. This phenomenon might be due to the substantial differences in the ionization states and gas distributions of the BLR among different AGNs, making it challenging to clearly reflect the effects of radiation pressure or accretion rate in the existing RM samples. On the other hand, the velocity-resolved ionization mapping (IM) technique proposed by Li et al. (2024) suggests that some observed velocity-resolved lags cannot be fully explained by simple inflow or outflow motions within a symmetric BLR, indicating a need for more complex geometric or kinematic configurations.
@@ 2.1 Lijiang Observations and Data Reduction
Data reduction was performed using the PyRAF111https://github.com/iraf-community/pyraf software, including bias subtraction, flat-fielding, wavelength calibration, spectral and photometric flux extraction, and flux calibration using spectrophotometric standard stars. The extraction aperture radius for both photometric and spectroscopic data was 424. For photometric data, we used several comparison stars near the target for relative flux calibration. For spectroscopic data, we obtained the fiducial spectrum of the comparison star during good weather conditions and used it for absolute flux calibration. Figure 2 presents the mean spectrum for each observing season. Note that our continuum measurements only utilize the -band data, while other photometric data will be used in future analyses of the accretion disk size.
@@ 2.2 Archival Time-domain Photometric Data
We also utilized publicly available photometric survey data from the All-Sky Automated Survey for Supernovae222https://www.astronomy.ohio-state.edu/asassn/index.shtml (ASAS-SN; Shappee et al., 2014; Hart et al., 2023) and the Zwicky Transient Facility333https://www.ztf.caltech.edu/ (ZTF; Masci et al., 2019) to investigate the long-term variability of NGC 4151 over the past decade. Both projects offer high sampling cadence and temporal overlap with our RM observations, enabling us to extend the time baseline of the continuum light curves and enhance the sampling cadence for measuring time lags.
@@
Since our -band wavelength coverage is closer to the -band of ASAS-SN and ZTF, we focused our analysis on the -band data while using the ASAS-SN -band to supplement data before 2017. For the long-term variability analysis, we primarily relied on the ASAS-SN -band data as a reference. RM measurements for the first year were directly obtained from Paper I, with ZTF -band light curves retrieved via Automatic Learning for the Rapid Classification of Events (ALeRCE) 444https://alerce.online/ (Förster et al., 2021; Sánchez-Sáez et al., 2021). In the following three observing seasons, due to the lower sampling rate of the ZTF data, we mainly used the ASAS-SN -band observations. Before analysis, data points with signal-to-noise ratio (S/N) below 5 were excluded, and observations were binned within 0.8-day intervals for each dataset.
@@ 3.1 Intercalibration of Photometric Data
The intercalibration was performed using the PyCALI555https://github.com/LiyrAstroph/pyCALI (Li et al., 2014) program, which employs a Bayesian approach assuming constant AGN component color and host-galaxy starlight during observations to compute flux differences between bands. This assumption is applicable to our data because: (1) the wavelength coverage of the , , and bands are close, so color variations can be ignored; (2) the Lijiang and ASAS-SN data utilize large fixed apertures for photometry, which can effectively mitigate the variation of host galaxy contribution with seeing; and (3) the ALeRCE data have already corrected the host galaxy contribution during image subtraction. The long-term light curve from ZTF is extracted using PSF photometry, where the contribution of an extended galaxy relative to the point source may vary with seeing. However, given a typical -band seeing of 21 (Bellm et al., 2019) and the relatively weak host galaxy of NGC 4151 compared to the AGN (see Chen et al., 2023, and Figure 2), the variation in the host galaxy contribution should only fluctuate within a small range in most cases.
@@ 3.2 Spectroscopic Measurements
Following the method outlined in Paper I, we employ spectral fitting to decompose the spectra, and the fitting strategy is essentially consistent across the observing seasons. Since Paper I completed the fitting analysis of the data from S1, we focus on the spectra from the subsequent three seasons in this work. In the new fitting process, we allow the power-law index to be a free parameter to better accommodate variations in the continuum shape. This is motivated by the fact that during the new observing period, NGC 4151 entered a bright state and exhibited dramatic variability. Such behavior can lead to non-negligible time lags between the continuum at different wavelengths (Zhou et al., 2024) and potential changes in the relative contribution of the Paschen continuum.
@@ 3.3 Lags, SMBH Masses, and Accretion Rates
The dimensionless accretion rate of the AGN is calculated using the following formula:
where denotes the mass accretion rates of the black hole, is the Eddington luminosity, , , and is the inclination angle of the accretion disk. We adopt an average inclination angle of cos = 0.75. The results show that during the four years of observations, the of NGC 4151 ranges from 0.01 to 0.23. Table 4 presents the broad emission-line widths and black hole masses, while the accretion rates are given in Table 3.
@@ 3.4 Examination of Relation in CL-AGN
Examining the change in the BLR size of AGN with luminosity is crucial for understanding the relation. We compared the H radius () and for each observing season with the relation established for objects with by Du et al. (2018b), which closely resembles the relation reported by Bentz et al. (2013). The results indicate that during our observations, NGC 4151 exhibited an “anti-breathing” phenomenon (Gilbert & Peterson, 2003; Goad et al., 2004; Cackett & Horne, 2006; Wang et al., 2020), where the time delay of the broad emission line decreased as the luminosity increased. Nevertheless, the data points from all four observations are basically located near the relation for AGNs with low .
Du et al. (2018b) found that when 3, the BLR lag can be shorter than the expected value from the relation, and this offset of lag is related to the accretion rate. Accordingly, we also examined the relationship between the relative offset of the H radius () and . We found that NGC 4151 is consistent with AGNs where , suggesting that the CL phenomenon may be associated with lower accretion rates. To verify that this is not a unique property of NGC 4151, we compiled an RM sample of CL-AGNs (Table 3) and conducted the same analysis. As shown in Figure 4, all CL-AGNs are sub-Eddington accretors and conform to the normal relation.
@@ 3.5 Velocity-resolved Signatures
Figure 5 shows the measurements of velocity-resolved lags for H, H, and H across each observing season. All three emission lines exhibit variability, with H and H showing more rapid changes. In contrast, H only demonstrates a different velocity-resolved lag pattern in the final year compared to the previous three years. Figure 6 presents the results of velocity-resolved BDs, which also exhibit changes across different observing seasons. These results are evident that the BLR geometry and kinematics of BLR in NGC 4151 are undergoing rapid changes on timescales of months to years.

### v2

@@ 1 Introduction
Over the past few decades, extensive observational and theoretical studies have revealed the basic picture of AGN structure: the central SMBH produces most of the observed continuum radiation through an accretion disk, which is surrounded by the broad-line region (BLR) composed of rapidly orbiting ionized gas emitting broad emission lines, a dust torus that can obscure the inner regions depending on the viewing angle, and the narrow-line region (NLR) extending to larger scales and emitting narrow emission lines (Antonucci, 1993; Urry & Padovani, 1995). Despite the huge success of the unified model in explaining the diverse observational features of AGNs, there remain many mysteries regarding the detailed physical processes and evolution of these structures, such as the observed accretion disk sizes often being larger than theoretically expected, the complex geometry and kinematics of the BLR, and the influence of radiation pressure on the internal physical properties of AGNs (e.g., Gaskell & Harrington, 2018; Pancoast et al., 2018; Villafaña et al., 2022; Chen et al., 2023; Liu et al., 2024; Li et al., 2024; Zhou et al., 2024).
@@
Additionally, some high-quality RM data allow for a detailed analysis of the emission line response at different line-of-sight velocities, enabling the inference of the geometry and kinematic properties of the BLR. This approach is known as the “velocity-resolved” RM method. These studies have revealed that, in addition to Keplerian motion, the BLR often exhibits complex kinematic features such as inflows or outflows (Denney et al., 2009; Bentz et al., 2010; Grier et al., 2013; Pancoast et al., 2014; Du et al., 2016; Lu et al., 2019; Li et al., 2021). Intriguingly, RM observations of the same AGN at different durations sometimes show significant changes in the kinematic features of the BLR (Grier et al., 2017; Lu et al., 2022; Chen et al., 2023; Yao et al., 2024). It is generally believed that changes in BLR kinematics may be related to variations in the accretion rate (Proga et al., 2000; Kollatschny & Zetzl, 2013), where higher accretion rates could lead to stronger radiation pressure driving more outflowing gas. However, Du et al. (2016) found no significant correlation between BLR kinematics and accretion rates through a sample of high accretion rate AGNs. This phenomenon might be due to the substantial differences in the ionization states and gas distributions of the BLR among different AGNs, making it challenging to clearly reflect the effects of radiation pressure or accretion rate in the existing RM samples. On the other hand, the velocity-resolved ionization mapping (IM) technique proposed by Li et al. (2024) suggests that some observed velocity-resolved lags cannot be fully explained by simple inflow or outflow motions within a symmetric BLR, indicating a need for more complex geometric or kinematic configurations.
@@ 2.1 Lijiang Observations and Data Reduction
Data reduction was performed using the PyRAF software (Science Software Branch at STScI, 2012), including bias subtraction, flat-fielding, wavelength calibration, spectral and photometric flux extraction, and flux calibration using spectrophotometric standard stars. The extraction aperture radius for both photometric and spectroscopic data was 424. For photometric data, we used several comparison stars near the target for relative flux calibration. For spectroscopic data, we obtained the fiducial spectrum of the comparison star during good weather conditions and used it for absolute flux calibration. Figure 2 presents the mean spectrum for each observing season. Note that our continuum measurements only utilize the -band data, while other photometric data will be used in future analyses of the accretion disk size.
@@ 2.2 Archival Time-domain Photometric Data
We also utilized publicly available photometric survey data from the All-Sky Automated Survey for Supernovae111https://www.astronomy.ohio-state.edu/asassn/index.shtml (ASAS-SN; Shappee et al., 2014; Hart et al., 2023) and the Zwicky Transient Facility222https://www.ztf.caltech.edu/ (ZTF; Masci et al., 2019) to investigate the long-term variability of NGC 4151 over the past decade. Both projects offer high sampling cadence and temporal overlap with our RM observations, enabling us to extend the time baseline of the continuum light curves and enhance the sampling cadence for measuring time lags.
@@
Since our -band wavelength coverage is closer to the -band of ASAS-SN and ZTF, we focused our analysis on the -band data while using the ASAS-SN -band to supplement data before 2017. For the long-term variability analysis, we primarily relied on the ASAS-SN -band data as a reference. RM measurements for the first year were directly obtained from Paper I, with ZTF -band light curves retrieved via Automatic Learning for the Rapid Classification of Events (ALeRCE) 333https://alerce.online/ (Förster et al., 2021; Sánchez-Sáez et al., 2021). In the following three observing seasons, due to the lower sampling rate of the ZTF data, we mainly used the ASAS-SN -band observations. Before analysis, data points with signal-to-noise ratio (S/N) below 5 were excluded, and observations were binned within 0.8-day intervals for each dataset.
@@ 3.1 Intercalibration of Photometric Data
The intercalibration was performed using the PyCALI (Li et al., 2014; Li, 2024) program, which employs a Bayesian approach assuming constant AGN component color and host-galaxy starlight during observations to compute flux differences between bands. This assumption is applicable to our data because: (1) the wavelength coverage of the , , and bands are close, so color variations can be ignored; (2) the Lijiang and ASAS-SN data utilize large fixed apertures for photometry, which can effectively mitigate the variation of host galaxy contribution with seeing; and (3) the ALeRCE data have already corrected the host galaxy contribution during image subtraction. The long-term light curve from ZTF is extracted using PSF photometry, where the contribution of an extended galaxy relative to the point source may vary with seeing. However, given a typical -band seeing of 21 (Bellm et al., 2019) and the relatively weak host galaxy of NGC 4151 compared to the AGN (see Chen et al., 2023, and Figure 2), the variation in the host galaxy contribution should only fluctuate within a small range in most cases.
@@ 3.2 Spectroscopic Measurements
Following the method outlined in Paper I, we employ spectral fitting using the DASpec software (Du, 2024) to decompose the spectra, and the fitting strategy is essentially consistent across the observing seasons. Since Paper I completed the fitting analysis of the data from S1, we focus on the spectra from the subsequent three seasons in this work. In the new fitting process, we allow the power-law index to be a free parameter to better accommodate variations in the continuum shape. This is motivated by the fact that during the new observing period, NGC 4151 entered a bright state and exhibited dramatic variability. Such behavior can lead to non-negligible time lags between the continuum at different wavelengths (Zhou et al., 2024) and potential changes in the relative contribution of the Paschen continuum.
@@ 3.3 Lags, SMBH Masses, and Accretion Rates
The dimensionless accretion rate of the AGN is calculated using the following formula(Du et al., 2018b):
where denotes the mass accretion rates of the black hole, is the Eddington luminosity, , , and is the inclination angle of the accretion disk. The inclination angle of NGC 4151, as derived from its NLR, is (cos ; Das et al., 2005; Fischer et al., 2013), which agrees well with the average inclination angle of (cos ; Nandra et al., 2007) found in AGNs using Fe K measurements. Given that the inclination angle is not measured for most AGNs, and to facilitate comparison with the results in Du et al. (2018b), we adopt the same average value of cos . The results show that during the four years of observations, the of NGC 4151 ranges from 0.01 to 0.23. Table 4 presents the broad emission-line widths and black hole masses, while the accretion rates are given in Table 3.
@@ 3.4 Examination of Relation in CL-AGN
Examining the change in the BLR size of AGN with luminosity is crucial for understanding the relation. We compared the H radius () and for each observing season with the relation established for objects with by Du et al. (2018b), which closely resembles the relation reported by Bentz et al. (2013). The results indicate that during our observations, NGC 4151 exhibited an “anti-breathing” phenomenon (Gilbert & Peterson, 2003; Goad et al., 2004; Cackett & Horne, 2006; Wang et al., 2020), where the time delay of the broad emission line decreased as the luminosity increased. The data points from the four observing seasons exhibit considerable scatter, though they do not deviate from the locations of AGNs with similar luminosity.
Du et al. (2018b) found that when 3, the BLR lag can be shorter than the expected value from the relation, and this offset of lag is related to the accretion rate. Accordingly, we also examined the relationship between the relative offset of the H radius () and . We found that NGC 4151 is consistent with AGNs where , suggesting that the CL phenomenon may be associated with lower accretion rates. To verify that this is not a unique property of NGC 4151, we compiled an RM sample of CL-AGNs (Table 3) and conducted the same analysis. As shown in Figure 7, all CL-AGNs are sub-Eddington accretors, and their distribution on the vs. plot is similar to that of AGNs with . This is also consistent with the fact that most CL-AGNs are classified as Population B sources in the Eigenvector 1 schema, as Population B is generally associated with lower accretion rates (Panda & Śniegowska, 2024).
@@ 3.5 Velocity-resolved Signatures
Figure 8 shows the measurements of velocity-resolved lags for H, H, and H across each observing season. All three emission lines exhibit variability, with H and H showing more rapid changes. In contrast, H only demonstrates a different velocity-resolved lag pattern in the final year compared to the previous three years. Figure 9 presents the results of velocity-resolved BDs, which also exhibit changes across different observing seasons. These results are evident that the geometry and kinematics of BLR in NGC 4151 are undergoing rapid changes on timescales of months to years.